# VideoRFT: Incentivizing Video Reasoning Capability in MLLMs via Reinforced Fine-Tuning

### v1

@@ 3.3 Overall Reward
  • Format Reward. During RL, we incorporate the widely-used format reward to guide the model in generating its reasoning process and final answer in a structured format. This reward, denoted as $R_{f}$, ensures that the model’s output adhere to a predefined structure: reasoning process must be enclosed within <think></think> tags, and answers within <answer></answer> tags. Compliance is verified via regular expression matching, and a binary reward is assigned accordingly.
@@
where $\mathbb{1}[R_{a}>0]$ is the indicator function that returns 1 if $R_{a}\!>\!0$ and 0 otherwise. It serves as a gate to control that the semantic-consistency reward $R_{s}$ is only activated when the accuracy reward $R_{a}$ is non-zero, which avoids reinforcing semantically plausible but factually incorrect reasoning.
@@ 4.1 Experimental Setup
Model Training. We follow the RFT to train VideoRFT in two stages: the warm-up SFT stage and the rule-based RL stage. Specifically, the SFT stage equips the model with the ability to generate correct responses for diverse questions, and is trained based VideoRFT-CoT-102K. The rule-based RL stage is based on VideoRFT-RL-310K using the reward in Eq. 6 to optimize structured reasoning and ensure factual validity.
Implementation Details. We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB per each. For efficiency, the video input is limited to 16 frames of resolution $128\times 28\times 28$ during training, where $28\!\times\!28$ is the size of each image patch, and $128$ denotes the number of patches. During inference, we increase the number of frames to $32$ and the resolution to $256\times 28\times 28$. For efficiency, we use a lightweight version of SigCLIP with 400M parameters in computing the semantic-consistency reward. The entire model is trained for one epoch of SFT followed by 1K steps of RL.
Table 1: Performance Comparison. The best results are highlighted in bold. : Results are obtained using larger input resolutions, up to $768\!\times\!28\!\times\!28$ and 768 sampled frames, while ours are $256\!\times\!28\!\times\!28$ and $32$.
@@
Model Pub VSI. VideoMMMU MMVU MV. TempCom. VideoMME
@@
$\bullet$ Open-Sourced Models
@@ 4.2 Main Result
As in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), open-sourced MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
Several key observations can be drawn from the results. First, compared to our base model, i.e., Qwen2.5-VL-7B, VideoRFT achieves significant improvements across all six benchmarks, e.g., +5.0% on VSI-Bench, +7.2% on MMVU, and +7.0 on VideoMME. This demonstrates the effectiveness of our approach in incentivizing video reasoning capabilities in MLLMs. Moreover, VideoRFT consistently outperforms all non-RL open-sourced MLLMs. Second, VideoRFT surpasses the proprietary GPT-4o on VSI-Bench by +2.8%, highlighting the strong potential of RFT in in bridging the performance gap with closed-source models in the task of video reasoning. Third, when compared to contemporaneous works, our model delivers the best overall performance, ranking first on four out of six benchmarks. This validates the superiority and generalization ability of our method in comparison with recent endeavors.
@@ 4.3 Diagnostic Experiment
Training Paradigm. To validate the effectiveness of the RFT training paradigm in our approach, we build two baseline approaches: i.e., SFT only and RL only. The former trains the model solely with supervised fine-tuning on VideoRFT-CoT-102K, while the latter, also known as the “zero” model in DeepSeek-R1, relies exclusively on RL without prior SFT. As seen from Table 2, the RL only surpasses the SFT only counterpart on four out of six datasets, indicating the capability of RL in stimulating more generalized reasoning capability. When combining both stages as in RFT, our full model VideoRFT achieves the best results, substantially outperforming the two baselines across all datasets. This highlights the complementary strengths of SFT for stable initialization and RL for reasoning enhancement in tackling video reasoning.
Reward Modeling. We further examine the effect of the reward defined in Eq. 6. Specifically, we compare two ablated variants: the first uses only uses the Format Reward and Accuracy Reward ($R\!=\!R_{f}\!+\!R_{a}$), while the second incorporates all three rewards directly ($R\!=\!R_{f}\!+\!R_{a}\!+\!R_{f}$). As seen, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness. Finally, our full reward formulation, which conditionally activates $R_{s}$ via gating, achieves the best overall results. Notably, it brings substantial gains especially for video reasoning benchmarks, i.e., +2.2% on VSI-Bench, +0.9% on VideoMMMU and +3.3% on MMVU.
@@ 5.1 Multimodal Reasoning in MLLMs
Enabling reasoning in MLLMs has become a central objective in recent research li2024enhancing ; AoTD ; MMCoT ; DDCoT . In the image domain, early works such as MMCoT MMCoT and DDCoT DDCoT disentangle perception and reasoning by treating visual understanding as input prompts for subsequent inference. In the video domain, VoT VoT and STEP STEP decompose video reasoning into predefined stages, employing template-based prompting to facilitate multi-step inference. DoraemonGPT DoraemonGPT models video understanding through symbolic memory and external tool sequences, yet still follows a modular reasoning paradigm. While these methods offer structured supervision, their rigid designs often limit generalization across diverse temporal and causal scenarios. Recently, rule-based RL have emerged as a promising paradigm for promoting multimodal reasoning in MLLMs. Pioneering efforts such as Visual-RFT VisualRFT , R1-VL R1VL , and Reason-RFT ReasonRFT directly adapt rule-based RL to image perception tasks. Follow-ups like Vision-R1 VisionR1 and R1-OneVision R1OneVision further demonstrate its effectiveness in enabling CoT reasoning on images. Concurrently, this paradigm has also been explored for video understanding VideoChatR1 ; VideoR1 ; TinyLLaVA-Video-R1 . Despite encouraging progress, these methods face a fundamental bottleneck: the lack of large-scale, high-quality video CoT datasets, which limits the full potential of RFT in the video domain. Our work addresses this gap by proposing a scalable and cognitively inspired pipeline to automatically mine high-quality CoT annotations for videos. Beyond this, we introduce a novel reward modeling strategy based on cross-modal semantic consistency, which explicitly guides MLLMs to generate visually grounded reasoning traces, which proves to be highly effective in improving model performance.
@@ 5.2 Multimodal CoT Dataset Construction
CoT has proven effective for enhancing the reasoning capabilities of LLMs by encouraging step-by-step reasoning CoT ; AutomaticCoT . Constructing high-quality CoT data in multimodal settings, particularly for video reasoning, remains a major challenge due to the temporal complexity and visual ambiguity of video data VideoEspresso ; VideoCoT ; VisualCoT . Recent works have explored CoT construction in both image and video domains. LLaVA-CoT llavacot , Vision-R1 VisionR1 , and R1-OneVision R1OneVision simply convert visual inputs to textual descriptions before reasoning. This often leads to hallucinations and weak semantic alignment. Video-R1 VideoR1 adopts a simplistic prompting strategy that encourages MLLMs to generate CoT by inserting “let me think", “wait". into responses. However, such CoTs merely mimic the surface form of human thinking without engaging in genuine reasoning. VideoEspresso VideoEspresso generates CoT data by prompting GPT-4o with a small set of selected key frames. Due to the sparse visual context and reliance on a text-only model, the generated CoTs often lack grounding in the actual video content and are prone to hallucinations. In contrast, our CoT data combines the reasoning ability of reasoning LLMs and the multimodal ability of MLLMs, ensuring the reasoning depth and visual grounding of CoT data. Moreover, we use cognition-inspired prompts to enable the reasoning model to generate CoT data that is more in line with human cognition.
@@ 6 Conclusion
In this work, we introduce VideoRFT, a novel approach for incentivizing cognitive video reasoning capabilities in MLLMs through reinforced fine-tuning. To accomplish this, we propose a cross-modal pipeline that generates high-quality cognitive video CoT data simulating human reasoning processes, resulting in two large-scale, high-quality datasets: VideoRFT-CoT-102K and VideoRFT-RL-310K. Furthermore, to strengthen the RL phase, we develop semantic-consistency guided reward modeling to explicitly encourage the alignment between reasoning traces and visual evidence. Extensive experiments across six video reasoning benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs. We expect this work lay a foundation for future efforts in RFT-based video reasoning.

### v3

@@ 3.3 Overall Reward
  • Format Reward. During RL, we incorporate the widely-used format reward to guide the model in generating its reasoning process and final answer in a structured format. This reward, denoted as $R_{f}$, ensures that the model’s output adheres to a predefined structure: the reasoning process must be enclosed within <think></think> tags, and answers within <answer></answer> tags. Compliance is verified via regular expression matching, and a binary reward is assigned accordingly.
@@
where $\mathbb{1}[R_{a}>0]$ is the indicator function that returns 1 if $R_{a}\!>\!0$ and 0 otherwise. It serves as a gate to control whether the semantic-consistency reward $R_{s}$ is only activated when the accuracy reward $R_{a}$ is non-zero, which avoids reinforcing semantically plausible but factually incorrect reasoning.
@@ 4.1 Experimental Setup
Model Training. We follow the RFT to train VideoRFT in two stages: the warm-up SFT stage and the rule-based RL stage. Specifically, the SFT stage equips the model with the ability to generate correct responses for diverse questions, and is trained based on VideoRFT-CoT-102K. The rule-based RL stage is based on VideoRFT-RL-310K using the reward in Eq. 6 to optimize structured reasoning and ensure factual validity. The RL training is implemented using the HuggingFace TRL library trl , and our codebase is built upon Open-R1 openr1 .
Implementation Details. We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB each. For efficiency, the video input is limited to 16 frames of resolution $128\times 28\times 28$ during training, where $28\!\times\!28$ is the size of each image patch, and $128$ denotes the number of patches. During inference, we increase the number of frames to $32$ and the resolution to $256\times 28\times 28$. For efficiency, we use a lightweight version of SigCLIP with 400M parameters in computing the semantic-consistency reward. The entire model is trained for one epoch of SFT followed by 1K steps of RL.
Table 1: Performance Comparison. The best results are highlighted in bold. : Results are obtained using larger input resolutions, up to $768\!\times\!28\!\times\!28$ and 768 sampled frames, while ours are $256\!\times\!28\!\times\!28$ and 3232.
@@
Model Pub VSI. VideoMMMU MMVU MV. TempC. VideoMME
@@
$\bullet$ Open-Source Models
@@ 4.2 Main Result
As in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), Open-Source MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
Several key observations can be drawn from the results. First, compared to our base model, i.e., Qwen2.5-VL-7B, VideoRFT achieves significant improvements across all six benchmarks, e.g., +5.0% on VSI-Bench, +7.2% on MMVU, and +7.0% on VideoMME. This demonstrates the effectiveness of our approach in incentivizing video reasoning capabilities in MLLMs. Moreover, VideoRFT consistently outperforms all non-RL Open-Source MLLMs. Second, VideoRFT surpasses the proprietary GPT-4o on VSI-Bench by +2.8%, highlighting the strong potential of RFT in bridging the performance gap with closed-source models in the task of video reasoning. Third, when compared to contemporaneous works, our model delivers the best overall performance, ranking first on four out of six benchmarks. This validates the superiority and generalization ability of our method in comparison with recent endeavors.
@@ 4.3 Diagnostic Experiment
Training Paradigm. To validate the effectiveness of the RFT training paradigm in our approach, we build two baseline approaches: i.e., SFT only and RL only. The former trains the model solely with supervised fine-tuning on VideoRFT-CoT-102K, while the latter, also known as the “zero” model in DeepSeek-R1, relies exclusively on RL without prior SFT. As seen from Table 2, the RL only surpasses the SFT only counterpart on four out of six datasets, indicating the capability of RL in stimulating more generalized reasoning capabilities. When combining both stages as in RFT, our full model VideoRFT achieves the best results, substantially outperforming the two baselines across all datasets. This highlights the complementary strengths of SFT for stable initialization and RL for reasoning enhancement in tackling video reasoning.
Reward Modeling. We further examine the effect of the reward defined in Eq. 6. Specifically, we compare two ablated variants: the first uses only the Format Reward and Accuracy Reward ($R\!=\!R_{f}\!+\!R_{a}$), while the second incorporates all three rewards directly ($R\!=\!R_{f}\!+\!R_{a}\!+\!R_{s}$). As seen, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness. Finally, our full reward formulation, which conditionally activates $R_{s}$ via gating, achieves the best overall results. Notably, it brings substantial gains especially for video reasoning benchmarks, i.e., +2.2% on VSI-Bench, +0.9% on VideoMMMU and +3.3% on MMVU.
@@ 5.1 Multimodal Reasoning in MLLMs
Enabling reasoning in MLLMs has become a central objective in recent research li2024enhancing ; AoTD ; MMCoT ; DDCoT . In the image domain, early works such as MMCoT MMCoT and DDCoT DDCoT disentangle perception and reasoning by treating visual understanding as input prompts for subsequent inference. In the video domain, VoT VoT and STEP STEP decompose video reasoning into predefined stages, employing template-based prompting to facilitate multi-step inference. DoraemonGPT DoraemonGPT models video understanding through symbolic memory and external tool sequences, yet still follows a modular reasoning paradigm. While these methods offer structured supervision, their rigid designs often limit generalization across diverse temporal and causal scenarios. Recently, rule-based RL has emerged as a promising paradigm for promoting multimodal reasoning in MLLMs. Pioneering efforts such as Visual-RFT VisualRFT , R1-VL R1VL , and Reason-RFT ReasonRFT directly adapt rule-based RL to image perception tasks. Follow-ups like Vision-R1 VisionR1 and R1-OneVision R1OneVision further demonstrate its effectiveness in enabling CoT reasoning on images. Concurrently, this paradigm has also been explored for video understanding VideoChatR1 ; VideoR1 ; TinyLLaVA-Video-R1 . Despite encouraging progress, these methods face a fundamental bottleneck: the lack of large-scale, high-quality video CoT datasets, which limits the full potential of RFT in the video domain. Our work addresses this gap by proposing a scalable and cognitively inspired pipeline to automatically mine high-quality CoT annotations for videos. Beyond this, we introduce a novel reward modeling strategy based on cross-modal semantic consistency, which explicitly guides MLLMs to generate visually grounded reasoning traces, and proves to be highly effective in improving model performance.
@@ 5.2 Multimodal CoT Dataset Construction
CoT has proven effective for enhancing the reasoning capabilities of LLMs by encouraging step-by-step reasoning CoT ; AutomaticCoT . Constructing high-quality CoT data in multimodal settings, particularly for video reasoning, remains a major challenge due to the temporal complexity and visual ambiguity of video data VideoEspresso ; VideoCoT ; VisualCoT . Recent works have explored CoT construction in both image and video domains. LLaVA-CoT llavacot , Vision-R1 VisionR1 , and R1-OneVision R1OneVision simply convert visual inputs to textual descriptions before reasoning. This often leads to hallucinations and weak semantic alignment. Video-R1 VideoR1 adopts a simplistic prompting strategy that encourages MLLMs to generate CoT by inserting “let me think”, “wait”, etc.into responses. However, such CoTs merely mimic the surface form of human thinking without engaging in genuine reasoning. VideoEspresso VideoEspresso generates CoT data by prompting GPT-4o with a small set of selected key frames. Due to the sparse visual context and reliance on a text-only model, the generated CoTs often lack grounding in the actual video content and are prone to hallucinations. In contrast, our CoT data combines the reasoning abilities of reasoning LLMs and the multimodal abilities of MLLMs, ensuring the reasoning depth and visual grounding of CoT data. Moreover, we use cognition-inspired prompts to enable the reasoning model to generate CoT data that is more in line with human cognition.
@@ 6 Conclusion
In this work, we introduce VideoRFT, a novel approach for incentivizing cognitive video reasoning capabilities in MLLMs through reinforced fine-tuning. To accomplish this, we propose a cross-modal pipeline that generates high-quality cognitive video CoT data simulating human reasoning processes, resulting in two large-scale, high-quality datasets: VideoRFT-CoT-102K and VideoRFT-RL-310K. Furthermore, to strengthen the RL phase, we develop semantic-consistency guided reward modeling to explicitly encourage the alignment between reasoning traces and visual evidence. Extensive experiments across six video reasoning benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs. We expect this work to lay a foundation for future efforts in RFT-based video reasoning.
Appendix A Additional Experiments
A.1 Effect of Model Scaling on Reasoning Ability
Table 3: Performance Comparison with Small-scale Models
  Video Reasoning Video Understanding
Model VSI. VideoMMMU MMVU MV. TempC. VideoMME
TinyLLaVA-Video-R1-3B - - 46.9 - 49.5 46.6
VideoRFT-3B 32.5 41.1 55.1 59.5 61.0 45.4
VideoRFT-7B 36.8 51.1 68.5 62.1 73.7 59.8
To further assess the scalability and robustness of our method under limited computational resources, we train an additional 3B variant of our model. While its performance is naturally lower than the original 7B version, the 3B model still performs competitively across benchmarks. Crucially, when compared to TinyLLaVA-Video-R1-3B, a contemporary 3B model specifically optimized for lightweight deployment, our 3B variant outperforms it, with gains of +8.2% on MMVU and +11.5% on TempCompass. This significant margin confirms the competitiveness of our approach even in small-scale settings, and underscores its strong generalization and reasoning abilities under resource-constrained conditions.
A.2 Hyperparameter Analysis
Table 4: Hyperparameter experiments for VideoRFT.
  Video Reasoning Video Understanding
Hyperparameter VSI-Bench VideoMMMU MMVU MVBench TempCompass VideoMME
$w=1$ 34.2 49.2 67.9 62.6 73.5 61.1
$w=2$ 36.8 51.1 68.5 62.1 73.7 59.8
$w=3$ 35.6 50.1 67.9 60.8 72.3 60.4
$w=4$ 35.4 49.7 67.6 62.0 73.1 59.1
To evaluate the effect of the scaling factor $w$ in the semantic reward, we conducted a hyperparameter study across multiple benchmarks. As shown in Table˜4, the performance varies with different values of $w$, with w=2w{=}2 consistently achieving the best or near-best results across all benchmarks. This suggests that w=2w{=}2 provides a favorable balance between semantic alignment and factual precision. In comparison, both smaller (w=1w{=}1) and larger (w=3w{=}3 or w=4w{=}4) values lead to a slight decline in performance, implying that insufficient or excessive emphasis on semantic similarity may hinder the overall reasoning and understanding capabilities. These findings highlight that while performance is relatively stable across a range of scaling values, setting w=2w{=}2 consistently yields optimal results, demonstrating the effectiveness of moderate semantic weighting.
Figure 8: Illustration of reasoning traces derived from VideoRFT in VideoMME.
A.3 Aha Moment in VideoRFT
Figure 8 illustrates additional instances of the “Aha Moment” behavior observed in VideoRFT. Prior to reaching a final decision, the model exhibits a human-like tendency to pause and scrutinize its intermediate reasoning steps. This reflective process, indicative of deliberative reasoning, is marked in red within the figure.
Appendix B Limitations and Future Directions
We consider this work a strong foundation for advancing video reasoning research using MLLMs. Several avenues warrant further investigation in future studies:
  • Handling of Challenging Video Scenarios. While our model demonstrates strong performance across various benchmark tasks, its effectiveness may be affected under complex conditions such as rapid motion or severe visual occlusion. Incorporating finer-grained visual signals and higher frame-rate sampling may help mitigate these challenges, enabling more accurate and robust visual representations.
  • CoT Data Reliance. The quality of the generated CoT annotations is closely tied to the capabilities of the underlying reasoning language model (e.g., DeepSeek-R1). Future work could explore leveraging more advanced reasoning models to further enhance the quality of CoT data, potentially leading to improved performance of VideoRFT.
Appendix C Potential Social Impacts
  • Positive Impacts. Enhanced video understanding enabled by VideoRFT can benefit various applications. In education, it facilitates the development of intelligent tutoring systems using video content. In security, it improves the efficiency and accuracy of surveillance video analysis. Moreover, it supports content moderation by aiding in the detection and filtering of inappropriate material.
  • Negative Impacts. There exists a risk of misuse. Misinterpretation of video content due to over-reliance on automated analysis could lead to the spread of misinformation. Additionally, in surveillance scenarios, the deployment of such systems may raise concerns regarding privacy and ethical use.
Appendix D Detailed Prompt Used in CoT Generation
Here, we provide detailed prompts for each step of the VideoRFT CoT generation process.
D.1 Video CoT Generation
[…]
For all videos, we sample them at 1FPS and input them into GPT-4o-mini, and generate a structured representation according to the prompt $P_{\text{rep}}$:
[…]
You are a video analysis assistant designed to produce rich, analytical per-frame captions from video inputs.
[…]
<Input Video>
[Uncaptioned image] </Input Video>
Task: 1. Overall Video Caption: video_caption: A concise 1-2 sentence (20-30 words) summary capturing the main theme or action of the video. 2. Per-Frame Metadata:
Uniformly sample frames and wrap them into a JSON list. Each element must include:
timestamp: "$H$:$M$:SS", aligned with the sampling interval. caption: 2-3 sentences (30-50 words) describing the scene in detail. key_elements: an object with fields: objects: list of detected objects/entities (strings) actions: list of ongoing actions or movements (strings) scene: e.g.“kitchen”, “urban street” notable_features: list of distinctive colors, textures, or patterns (strings) spatial_relations: list of spatial relationships (e.g.“cup on table”, “person left of car”) human_attributes: object or null. If present, include: * gender: “male”, “female”, “unknown” * clothing: brief description * posture: “standing”, “sitting” potential_interactions: list of possible interactions (strings)
General Instructions: Temporal Consistency: Reference continuing actions from the previous frame and highlight any changes. Uncertainty: If confidence<<0.6 or object visibility<<50%, append [Uncertain] to the caption. Implied Actions: Describe preparatory movements (e.g.“hand reaching toward door handle” vs. “holding door handle”). Output Requirements: Wrap all per-frame objects into a single JSON list matching the number of sampled frames.
[…]
We invoke DeepSeek-R1 to answer the question and extracts its step-by-step reasoning outputs with the prompt $P_{\text{cog}}$ as the initial CoT.
[…]
You are an AI assistant helping a user answer questions about a video. When the user asks a question, you respond by imagining you are watching the video with full attention, just like a human would. Your task is to reason visually and logically about the video content to answer the user’s question.
[…]
Video content: <Overall Video Caption>
Frame-level metadata: <Per-Frame Metadata> Question: <Question> Please think about this question as if you were a human pondering deeply. It’s encouraged to include self-reflection or verification in the reasoning process. <Corresponding Answer Format Template>
The <Corresponding Answer Format Template> in the prompt is dynamically selected from the following templates based on the question type:
Table 5: Answer Format Templates for Different Question Types
Question Type Template
Multiple Choice Please provide only the single option letter (e.g., A, B, C, D, etc.) within the <answer> </answer> tags.
Numerical Please provide the numerical value within the <answer> </answer> tags.
OCR Please transcribe text from the image/video clearly and provide your text answer within the <answer> </answer> tags.
Free-form Please provide your text answer within the <answer> </answer> tags.
Regression Please provide the numerical value within the <answer> </answer> tags.
[…]
We employ a cross-modal refinement process to ensure the CoT aligns with the video content, using the prompt $P_{\text{cross}}$:
[…]
You are a multimodal reasoning expert. Your task is to revise hallucinations and errors in the chain-of-thought (CoT) based on the visual content of the provided video. Do not significantly alter the original CoT logic or content, and ensure the final conclusion remains the same.
[…]
<Input Video>
[Uncaptioned image] </Input Video>
Question: <Question> Original CoT: <Original CoT> Output format:
Strictly follow this format. Return only the revised CoT and no additional explanation:
<think>[Revised CoT]</think>
D.2 Image CoT Generation
In addition to the main video data in the VideoRFT-CoT-102K, we have also designed specific prompts specifically for image data, and used the similarly CoT generation process to generate CoT.
[…]
Due to the differences among different image datasets (the task focuses of the datasets are different), we designed different prompts for each image dataset in structured image representation phase.
Common System Prompt for All Image Representations You are a vision-language expert. Your task is to analyze the provided image and output a detailed, modular description in JSON format.
The following prompts are used for different image datasets, all sharing the above system prompt:
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given a natural photograph. Output only valid JSON with two keys: 1. Overall Image Caption: a 2-3 sentence narrative describing the scene, objects, actions, and context. 2. Image Metadata: an object containing: objects: list of objects with id, type, color, size, and bbox text: list of text elements with content and bbox scene_context: environment and activity relations: list of subject-predicate-object relations
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given a synthetic 3D scene. Output only valid JSON with: 1. Overall Image Caption: a brief summary of number of objects and overall layout. 2. Image Metadata: objects: list of objects with id, shape, color, size, material, and coordinates spatial_relations: list of spatial relations between objects
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given a line-drawing or diagram. Output only valid JSON with: 1. Overall Image Caption: a concise paragraph describing the diagram’s purpose. 2. Image Metadata: primitives: list of geometric primitives with type, label, and bbox annotations: list of relations between primitives measurements: list of measurements with primitive IDs and values
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given an image containing printed or handwritten text. Output only valid JSON with: 1. Overall Image Caption: one sentence summarizing the text context. 2. Image Metadata: text_items: list of text elements with content, bbox, and style
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given a multi-panel scientific figure. Output only valid JSON with: 1. Overall Image Caption: a paragraph overviewing the figure’s subject. 2. Image Metadata: panels: list of panels with elements and process arrows
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given a chart image. Output only valid JSON with: 1. Overall Image Caption: a 1-2 sentence summary of chart type and key trend. 2. Image Metadata: chart_type: type of chart axes: list of axis information series: list of data series legend: list of legend entries
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given a textbook problem image. Output only valid JSON with: 1. Overall Image Caption: a summary of the problem context. 2. Image Metadata: equations: list of equations with LaTeX and bbox diagram_parts: list of diagram elements givens: list of given values
[…]
<Input Image>
[Uncaptioned image] </Input Image>
You are given an indoor scene or floorplan image. Output only valid JSON with: 1. Overall Image Caption: a 2-3 sentence narrative of the space. 2. Image Metadata: rooms_or_sections: list of room information furniture: list of furniture items with position and orientation annotations: list of structural elements
[…]
We invoke DeepSeek-R1 to answer the question and extracts its step-by-step reasoning outputs with the prompt $P_{\text{cog}}$ as the initial CoT.
[…]
You are an AI assistant helping a user answer questions about an image. When the user asks a question, you respond by imagining you are looking at the image with full attention, just like a human would. Your task is to reason visually and logically about the image content to answer the user’s question.
[…]
Image content: <Overall Image Caption>
Image metadata: <Image Metadata> Question: <Question> Please think about this question as if you were a human pondering deeply. It’s encouraged to include self-reflection or verification in the reasoning process. <Corresponding Answer Format Template>
[…]
We employ a cross-modal refinement process to ensure the CoT aligns with the image content, using the prompt $P_{\text{cross}}$:
[…]
You are a multimodal reasoning expert. Your task is to revise hallucinations and errors in the chain-of-thought (CoT) based on the provided image. Do not significantly alter the original CoT logic or content, and ensure the final conclusion remains the same.
[…]
<Input Image>
[Uncaptioned image] </Input Image>
Question: <Question> Original CoT: <Original CoT> Output format:
Strictly follow this format. Return only the revised CoT and no additional explanation:
<think>[Revised CoT]</think>